An athle run with speed of 12m/s .Determine the longest jump he can take. Solve with complete explanation.


1
Expert's answer
2021-05-23T18:54:58-0400

"l=\frac{v^2 \sin 2\alpha}{g},"

"l=l_{max}~\text {if }\sin2\alpha=1~(\alpha=45°),"

"l_{max}=\frac{v^2}{g}=\frac{12^2}{9.8}=14.7~\frac ms."